# On Adversarial Attacks In Acoustic Drone Localization

### v1

@@ B.1 Modeling Noise
Since we are the first work to address adversarial attacks over acoustic drone localization, in our experiments we wish to assess the impact of acoustic adversarial attacks over drone localization with minimal assumptions over the attacker and the environment. Therefore, in all previous experiments we assume an idealized environment where the only sound sources are the drone and the adversarial perturbation. This assumption maximizes the attacker’s flexibility and isolates external interference. Nonetheless, in order to provide an initial assessment of our findings in into real-world conditions, in this section we test the potential of our developed attacked under the presence of white noise.
@@
B.2 Targeted Attacks
@@
A central motivation for the study of adversarial attacks in drone localization is the possibility of a potential attacker directing the drone to a specific location, where it could potentially pose harm to the original user or third-party individuals. We therefore also analyze the affect of targeted adversarial attacks, brought forth in Figure 7. In this figure we show spatial error distributions for targeted attacks in different locations (marked in green on the heatmap, and textually on each sub-figure caption). Note that unlike heatmaps for non-targeted attacks, here the error reported on the heatmap is RMS from adversarially desired target (so an all-black heatmap means perfectly successful attack).
@@
We observe that the success of targeted attacks depends on target location. Subfigures a-c present very successful adversarial targeted attack, where the adversarially-desired target is predicted almost everywhere in the map. Figures g-i present the complementary scenario, where the clean and attacked error maps are nearly identical for all bounds. We speculate this difference stems from the fact that the attack targeted at $[-2,-4.5]$ (namely, the ”successful” attack) is focused on a more-likely low-certainty area (given the room shape in the area of $[-2,0]$, that is non-representative of the common shape at other scene regions), contrary to the attack at $[8,-4.5]$ (an area where the attack allegedly failed).
@@
B.3 Fixed-perturbation source RMS distribtution
Figure 8 presents results similar to figure 5 along with fixed-source location attack optimization. Results support our findings from section 4.2, stating that location optimization holds limited contribution for attack success in our tested setting. Similar conclusions can be drawn from results over the office (Figure 9) and room Figure 10 environments, also complementing the partial results for location optimization in the two latter environments, as shown in Figure 5.
@@ B.3 Fixed-perturbation source RMS distribtution
B.4 Perturbation Delineation Uncertainty
Our perturbation recovery algorithm recovers the exact perturbation waveform up to uncertainty in the sampled perturbation value at moment $t=0$, in this section we further evaluate our method by assessing the affect of this deviation over model performance. Namely, for a clean input signal $s_{drone}(t)$, our recovered signal in Section 3.4 is $(s_{drone}+\delta)(t)$ for some unknown scalar offset $\delta\in\mathbf{R}$ (that equals the perturbation response at the microphone at moment $t=0$). Our primary approach for evaluating the repercussion of uncertainty in $\delta$ is by estimation of the spatial distribution of location-wise perturbation values at $t=0$ (denoted $s_{\mathrm{p}}^{\prime}(t=0;\mu)$ and location gradient infinity norms across different sensor locations $\mu$ $\mathcal{L}_{\infty}(\frac{\partial\mathcal{F}(s_{drone})}{\partial s_{drone}})$.
Figure 13: Spatial distribution of maximum location prediction response to change in $\delta$ demonstrating responses typically several magnitudes of order lower compared to absolute locations
@@ C.1 Clean Model Training
For training our ”clean” localization model (to later be attacked) we train both the forward model and the transformer-encoder inverse model from (Shor et al., 2024)
@@ C.1.1 Forward Model
We parameterize the self-sound of each rotor with 16 point sound-sources, and optimize using L-BFGS for 50 iterations. The self-sound is then used along with the scene-specific pre-trained RIRs published in (Luo et al., 2022) for each test environment.
@@ C.1.2 Localization Model
We train the localization model with dataset generated by sampling each environment with a $5\times 5$ cm density. For each such drone location we computed the sound sampled by the microphone array in 32 azimuthal orientations uniformly spaced within the $[0,2\pi]$ range. With this dataset of sounds sampled by the microphones at different drone locations and orientations ( 125000 samples for the apartment, 65000 for office and 42000 for room) we train the localization model, using the transformer-encoder architecture from (Shor et al., 2024). We subsample each waveform from a temporal dimension of  12000 timesteps to 1200 using a learned linear operator. This subsampled input is then fed into a 3-layered transformer-encoder model with hidden dimension of 1024. This model is trained on batches of 128 for 80 epochs on an NVIDIA RTX-2080 GPU, with phase modulation optimization of 10 sine basis components. Model parameters are optimized with an Adam optimizer, using a learn-rate of 0.00001 for the localization model and a learn rate of 0.05 for the phase modulation parameters.
@@ C.2 Perturbation Optimization
0:  Perturbed waveform $s_{\mu}$, period time $T$
0:  Recovered sampled perturbation waveform $s_{\mathrm{r}}$
1:  Initialize $s_{\mathrm{r}}=[\ ]$ Recovered Waveform
2:  for $j=1$ to $T$ do
3:     sample $s_{\mu}(t;j)$ Modulate by $j$ timesteps
4:     update $s_{\mathrm{r}}(t=j)=s_{\mu}(t;j)-s_{\mu}(t;0)$ Calculate perturbation at timestep $j$
5:  end for
6:  return $s_{\mathrm{r}}$
@@ Appendix D Resource Analysis
We denote that the batch-size of 128 mentioned in section C is used by internally batching RIR computations to batches of 2, and only inserting the batch-size of 128 onto the localization model.

### v2

@@ 4.4 Noise Modeling
Since we are the first work to address adversarial attacks over acoustic drone localization, in our experiments we wish to assess the impact of acoustic adversarial attacks over drone localization with minimal assumptions over the attacker and the environment. Therefore, in all previous experiments we assume the only sound sources are the drone and the adversarial perturbation. This assumption maximizes the attacker’s flexibility and isolates external interference. Nonetheless, in order to provide an initial assessment of our findings in real-world conditions, in this section we test the potential of our developed attacked under the presence of white noise.
@@
5 Discussion
In this work, we presented a first study of adversarial attacks on acoustic localization. We formulated a framework for generating universal acoustic perturbations and analyzed their impact on localization accuracy under various conditions. We also introduced a novel algorithm for recovering the adversarial signal. Our attacks increased localization RMS error by up to 37%. Our approach also demonstrated that adversarial optimization of source location has minimal impact on attack effectiveness, enabling efficient attacks with lower computational cost. As a by-product, we trained and evaluated a self-sound-reliant drone localization model on real acoustic data, expanding prior work that focused solely on simulations.
While our work provides a foundational framework for adversarial attacks and defenses in acoustic localization, several limitations remain. First, in this study we focus on 2D acoustic localization. Many real-world localization scenarios incorporate a 6-DoF regression problem, calling for an extension of our contributions to higher-dimension localization. Second, in order to isolate the marginal impact of the adversarial attack, our work assumes a simplistic noise model, that does not grasp the chaotic acoustic nature found in some scenes. Furthermore, while efficient, our perturbation delineation algorithm must be applied for an entire drone cycle for every moment in time $[0,T]$, which requires several seconds to delineate a single perturbative waveform.
Another promising avenue for future research would be the exploration of more-complex adversarial settings, that are not discussed in this work. While our experimental evaluation is focused on a single stationary perturbation source, the proposed defense is theoretically extensible to more-general scenarios. Since acoustic signals in a linear medium superimpose, a multi-source attack effectively manifests as a single complex waveform at the sensor array. Because our defense relies on the independence of external sound sources from the drone’s internal phase modulation mechanism, it is potentially capable of extending to any number of stationary, non-adaptive sources without modification. Another viable adversarial setting for future exploration would be active real-time adversarial adaptation done by the attacker. In the case of our proposed defense, it is constrained by causality. The propagation delay of sound and the processing time required to estimate the drone’s phase prevent the attacker from reacting instantaneously to the defense mechanism. We hope our findings help pave the way for the development of robust and secure acoustic perception in future autonomous systems.
Appendix A Supplementary Layout
In the following sections we report additional details complementing the main paper, ordered as follows - 1) Section B includes supplementary results for the three main types of experiments conducted in the paper -
we note that our scatter-plot heatmap figures for the room_2 and office_4 environments may depict environment shapes slightly different than those known from Chang et al. (2017); Straub et al. (2019) environments. The reason is we do not evaluate localization in locations where the drone cannot be present, due to its simulated physical dimensions from Serussi et al. (2024).
2) Section C provides in-depth implementation, optimization and training-regime details. 3) In Section D we discuss the added computational resources required in adversarial source location optimization, to further stress the importance of our findings from section 4.2, where we claim source-location optimization could prove cost-ineffective in some cases.
4) Section E further provide added details for our proposed perturbation delineation algorithm.
Appendix B Additional Results
B.1 Targeted Attacks
@@
A central motivation for the study of adversarial attacks in drone localization is the possibility of a potential attacker directing the drone to a specific location, where it could potentially pose harm to the original user or third-party individuals. We therefore also analyze the effect of targeted adversarial attacks, brought forth in Figure 7. In this figure we show spatial error distributions for targeted attacks in different locations (marked in green on the heatmap, and textually on each sub-figure caption). Note that unlike heatmaps for non-targeted attacks, here the error reported on the heatmap is RMS from adversarially desired target (so an all-black heatmap means perfectly successful attack).
@@
We observe that the success of targeted attacks depends on target location. Subfigures a-c present very successful adversarial targeted attack, where the adversarially-desired target is predicted almost everywhere in the map. Figures g-i present the complementary scenario, where the clean and attacked error maps are nearly identical for all bounds. We speculate this difference stems from the fact that the attack targeted at [2,4.5][-2,-4.5] (namely, the "successful" attack) is focused on a more-likely low-certainty area (given the room shape in the area of [2,0][-2,0], that is non-representative of the common shape at other scene regions), contrary to the attack at [8,4.5][8,-4.5] (an area where the attack allegedly failed).
@@
B.2 Fixed-perturbation source RMS distribution
Figure 8 presents results similar to Figure 5 along with fixed-source location attack optimization. Results support our findings from Section 4.2, stating that location optimization holds limited contribution for attack success in our tested setting. Similar conclusions can be drawn from results over the office (Figure 9) and room (Figure 10) environments, also complementing the partial results for location optimization in the two latter environments, as shown in Figure 5.
@@ B.2 Fixed-perturbation source RMS distribution
B.3 Perturbation Delineation Uncertainty
Our perturbation recovery algorithm recovers the exact perturbation waveform up to uncertainty in the sampled perturbation value at moment $t=0$, in this section we further evaluate our method by assessing the effect of this deviation over model performance. Namely, for a clean input signal $s_{drone}(t)$, our recovered signal in Section 3.4 is $(s_{drone}+\delta)(t)$ for some unknown scalar offset $\delta\in\mathbf{R}$ (that equals the perturbation response at the microphone at moment $t=0$). Our primary approach for evaluating the repercussion of uncertainty in $\delta$ is by estimation of the spatial distribution of location-wise perturbation values at $t=0$ (denoted $s_{\mathrm{p}}^{\prime}(t=0;\mu)$ and location gradient infinity norms across different sensor locations $\mu$ - $\mathcal{L}_{\infty}(\frac{\partial\mathcal{F}(s_{drone})}{\partial s_{drone}})$.
Figure 13: Spatial distribution of maximum location prediction response to change in $\delta$ - demonstrating responses typically several magnitudes of order lower compared to absolute locations
@@ C.1 Clean Model Training
For training our "clean" localization model (to later be attacked) we train both the forward model and the transformer-encoder inverse model from Serussi et al. (2024)
@@ C.1.1 Forward Model
We parameterize the self-sound of each rotor with 16 point sound-sources, and optimize using L-BFGS for 50 iterations. The self-sound is then used along with the scene-specific pre-trained RIRs published in Luo et al. (2022) for each test environment.
@@ C.1.2 Localization Model
We train the localization model with dataset generated by sampling each environment with a $5\times 5$ cm density. For each such drone location we computed the sound sampled by the microphone array in 32 azimuthal orientations uniformly spaced within the $0,2\pi]$ range. With this dataset of sounds sampled by the microphones at different drone locations and orientations ( 125000 samples for the apartment, 65000 for office and 42000 for room) we train the localization model, using the transformer-encoder architecture from Serussi et al. (2024). We subsample each waveform from a temporal dimension of  12000 timesteps to 1200 using a learned linear operator. This subsampled input is then fed into a 3-layered transformer-encoder model with hidden dimension of 1024. This model is trained on batches of 128 for 80 epochs on an NVIDIA RTX-2080 GPU, with phase modulation optimization of 10 sine basis components. Model parameters are optimized with an Adam optimizer, using a learn-rate of 0.00001 for the localization model and a learn rate of 0.05 for the phase modulation parameters.
@@ Appendix D Resource Analysis
we note that the batch-size of 128 mentioned in Section C is used by internally batching RIR computations to batches of 2, and only inserting the batch-size of 128 onto the localization model.